# Muon Optimizes Under Spectral Norm Constraints

### v1

@@ Appendix B Deferred proofs
$\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})-\nabla\mathcal{F}(\mathbf{X}_{t-1})\right\rVert_{\textup{F}}$ $\leq L\left\lVert\mathbf{X}_{t}-\mathbf{X}_{t-1}\right\rVert_{\textup{F}}=\eta L\left\lVert\operatorname{msgn}(\widetilde{\mathbf{M}}_{t})-\lambda\mathbf{X}_{t}\right\rVert_{\textup{F}}$ (30)
$\leq\eta C_{\mathcal{K}}L\left\lVert\operatorname{msgn}(\widetilde{\mathbf{M}}_{t})-\lambda\mathbf{X}_{t}\right\rVert_{\textup{op}}\leq 2\eta C_{\mathcal{K}}L,$
@@
$\leq 2\eta C_{\mathcal{K}}L\sum_{k=1}^{t}\beta_{2}^{t-k}+\beta_{2}^{t}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}}\leq\frac{2\eta C_{\mathcal{K}}L}{1-\beta_{2}}+\beta_{2}^{t}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}},$
where the third line iterates and expands the first two lines, the fourth line uses the triangle inequality, the fifth line uses (30) and
$\sum_{k=1}^{t}\beta_{2}^{t-k}\leq\sum_{j=0}^{\infty}\beta_{2}^{j}=\frac{1}{1-\beta_{2}}.$
Thus
$\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{F}}$ $=\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\beta_{1}\mathbf{M}_{t-1}-(1-\beta_{1})\nabla\mathcal{F}(\mathbf{X}_{t-1})\right\rVert_{\textup{F}}$
@@
$\leq 2\eta C_{\mathcal{K}}L+\beta_{1}\left(\frac{2\eta C_{\mathcal{K}}L}{1-\beta_{2}}+\beta_{2}^{t-1}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}}\right)$
$=\frac{2\eta C_{\mathcal{K}}L(1+\beta_{1}-\beta_{2})}{1-\beta_{2}}+\beta_{1}\beta_{2}^{t-1}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}},$
where the third line uses the triangle inequality and the fourth line uses (30) and (31). ∎
@@
$\widehat{\mathcal{V}}_{t}:=\mathcal{V}(\mathbf{X}_{t})-\Gamma_{t},\quad\text{where }\Gamma_{t}:=\sum_{s=0}^{t-1}\gamma_{s}.$
@@
[…]
By (32), the nonnegativity of $\mathcal{V}$, and $\lim_{t\to\infty}\Gamma_{t}<\infty$, it follows that $\{\widehat{\mathcal{V}}_{t}\}_{t\in\mathbb{N}}$ is a supermartingale with $\sup_{t\in\mathbb{N}}\mathbb{E}\left[\widehat{\mathcal{V}}(\mathbf{X}_{t})^{-}\right]<\infty$. By Doob’s supermartingale convergence theorem, we conclude that $\{\widehat{\mathcal{V}}_{t}\}_{t\in\mathbb{N}}$ and hence $\{\mathcal{V}(\mathbf{X}_{t})\}_{t\in\mathbb{N}}$ converges almost surely to a random variable with finite expectation.
@@
$\sum_{t=0}^{\infty}\mathbb{E}[\alpha_{t}h(\mathbf{X}_{t+\ell})]=\sum_{t=\ell}^{\infty}\alpha_{t}\mathbb{E}[h(\mathbf{X}_{t})]<\infty.$
@@
$00\leq\mathbb{E}\left[\liminf_{t\to\infty}h(\mathbf{X}_{t})\right]\leq\liminf_{t\to\infty}\mathbb{E}[h(\mathbf{X}_{t})]=$ 0
@@
[…]
by Lemma 8. By the lower semicontinuity and nonnegativity of $h$ and the (almost sure) boundedness of $\mathbf{X}_{t}$, the $\omega$-limit set $\mathcal{M}$ of $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$ must satisfy $h(\mathbf{X})=0$ for all $\mathbf{X}\in\mathcal{M}$, i.e.
$\mathcal{M}\subseteq\{\mathbf{X}\in\mathbb{X}\mid h(\mathbf{X})=0\}.$

### v2

@@ Appendix B Deferred proofs
$\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})-\nabla\mathcal{F}(\mathbf{X}_{t-1})\right\rVert_{\textup{F}}$ $\leq L\left\lVert\mathbf{X}_{t}-\mathbf{X}_{t-1}\right\rVert_{\textup{F}}=\eta_{t-1}L\left\lVert\operatorname{msgn}(\widetilde{\mathbf{M}}_{t})-\lambda\mathbf{X}_{t}\right\rVert_{\textup{F}}$ (29)
$\leq\eta_{t-1}C_{\mathcal{K}}L\left\lVert\operatorname{msgn}(\widetilde{\mathbf{M}}_{t})-\lambda\mathbf{X}_{t}\right\rVert_{\textup{op}}\leq 2\eta_{t-1}C_{\mathcal{K}}L,$
@@
$\leq 2C_{\mathcal{K}}L\sum_{k=1}^{t}\beta_{2}^{t-k}\eta_{k-1}+\beta_{2}^{t}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}},$
where the third line iterates and expands the first two lines, the fourth line uses the triangle inequality, and the fifth line uses (29). Thus
$\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\widetilde{\mathbf{M}}_{t}\right\rVert_{\textup{F}}$ $=\left\lVert\nabla\mathcal{F}(\mathbf{X}_{t})+\beta_{1}\mathbf{M}_{t-1}-(1-\beta_{1})\nabla\mathcal{F}(\mathbf{X}_{t-1})\right\rVert_{\textup{F}}$ (31)
@@
$\leq 2\eta_{t-1}C_{\mathcal{K}}L+\beta_{1}\left(2C_{\mathcal{K}}L\sum_{k=1}^{t-1}\beta_{2}^{t-k-1}\eta_{k-1}+\beta_{2}^{t-1}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}}\right)$
$=2\eta_{t-1}C_{\mathcal{K}}L+2\beta_{1}C_{\mathcal{K}}L\sum_{k=1}^{t-1}\beta_{2}^{t-k-1}\eta_{k-1}+\beta_{1}\beta_{2}^{t-1}\left\lVert\nabla\mathcal{F}(\mathbf{X}_{0})+\mathbf{M}_{0}\right\rVert_{\textup{F}},$
where the third line uses the triangle inequality and the fourth line uses (29) and (30). The result follows upon setting $\eta_{t}=\eta$ and using $\sum_{k=1}^{t-1}\beta_{2}^{t-k-1}\leq\sum_{j=0}^{\infty}\beta_{2}^{j}=\frac{1}{1-\beta_{2}}.$
@@
$\widehat{\mathcal{V}}_{t}:=\mathcal{V}(\mathbf{X}_{t})-g_{t},\quad\text{where }g_{t}:=\sum_{s=0}^{t-1}\gamma_{s}.$
@@
$\mathbb{E}\left[\widehat{\mathcal{V}}_{t+1}\mid\mathcal{F}_{t}\right]-\widehat{\mathcal{V}}_{t}=\mathbb{E}[\mathcal{V}(\mathbf{X}_{t+1})\mid\mathcal{F}_{t}]-\mathcal{V}(\mathbf{X}_{t})-\gamma_{t}\leq-\alpha_{t}h(\mathbf{X}_{t+\ell})\leq 0\text{ a.s.}$ (32)
By (32), the nonnegativity of $\mathcal{V}$, and $\lim_{t\to\infty}g_{t}<\infty$, it follows that $\{\widehat{\mathcal{V}}_{t}\}_{t\in\mathbb{N}}$ is a supermartingale with $\sup_{t\in\mathbb{N}}\mathbb{E}\left[\widehat{\mathcal{V}}(\mathbf{X}_{t})^{-}\right]<\infty$. By Doob’s supermartingale convergence theorem, we conclude that $\{\widehat{\mathcal{V}}_{t}\}_{t\in\mathbb{N}}$ and hence $\{\mathcal{V}(\mathbf{X}_{t})\}_{t\in\mathbb{N}}$ converges almost surely to a random variable with finite expectation.
@@
$\sum_{t=0}^{\infty}\mathbb{E}[\alpha_{t}h(\mathbf{X}_{t+\ell})]=\sum_{t=\ell}^{\infty}\alpha_{t-\ell}\mathbb{E}[h(\mathbf{X}_{t})]<\infty.$
@@
$0\leq\mathbb{E}\left[\liminf_{t\to\infty}h(\mathbf{X}_{t})\right]\leq\liminf_{t\to\infty}\mathbb{E}[h(\mathbf{X}_{t})]=0$
@@
$\liminf_{t\to\infty}h(\mathbf{X}_{t})=0\text{ a.s.}$
by Lemma 8. By the lower semicontinuity and nonnegativity of $h$ and the (almost sure) boundedness of $\mathbf{X}_{t}$, the $\omega$-limit set $\mathbb{M}$ of $\{\mathbf{X}_{t}\}_{t\in\mathbb{N}}$ must satisfy $h(\mathbf{X})=0$ for all $\mathbf{X}\in\mathbb{M}$, i.e.
$\mathbb{M}\subseteq\{\mathbf{X}\in\mathbb{X}\mid h(\mathbf{X})=0\}.$